OpenStudy (anonymous):

A rain gutter is to be constructed of aluminum sheets 12 inches wide. After marking off a length of 4 inches from each edge, this length is bent up at an angle Θ. The area A of the opening may be expressed as the function: A(Θ) = 16 sin Θ • (cos Θ + 1). If Θ = 90°, what is the area of the opening?

OpenStudy (anonymous):

I got 16.0 in^2

OpenStudy (anonymous):

Yup since \[A(90)=16sin(90)*(cos(90)+1)\] \[A(90)=16(1)*(0+1)=16\]

OpenStudy (anonymous):

yay